Assignment Status Click here for detailed view Pro
Answer #1

(a)

moment of inertia

I = M*R^2

(b)

torque on yo-yo = I * alpha

(0.5R * T) = M*R^2 *( a / 0.5R)

where T = tension in string

T = 4M*a

Fnet = m*a in vertical,

2M*g - T = 2M*a

2M*g - 4M*a = 2M*a

a = g / 3

a = 9.81 / 3 = 3.27 m/s^2

using equation of motion

v^2 - u^2 = 2*a*s

v^2 - 0 = 2*3.27*H

V = 2.56 * sqrt(H)

(c)

H = 0.43 m

v = 2.56 * sqrt(0.43)

v = 1.67 m/s

answer